### v2

# Mozart’s Touch: A Lightweight Multi-modal Music Generation Framework Based on Pre-Trained Large Models

Abstract.
In recent years, AI-Generated Content (AIGC) has witnessed rapid advancements, facilitating the generation of music, images, and other forms of artistic expression across various industries. However, researches on general multi-modal music generation model remain scarce. To fill this gap, we propose a multi-modal music generation framework Mozart’s Touch. It could generate aligned music with the cross-modality inputs, such as images, videos and text. Mozart’s Touch is composed of three main components: Multi-modal Captioning Module, Large Language Model (LLM) Understanding & Bridging Module, and Music Generation Module. Unlike traditional approaches, Mozart’s Touch requires no training or fine-tuning pre-trained models, offering efficiency and transparency through clear, interpretable prompts. We also introduce ”LLM-Bridge” method to resolve the heterogeneous representation problems between descriptive texts of different modalities. We conduct a series of objective and subjective evaluations on the proposed model, and results indicate that our model surpasses the performance of current state-of-the-art models. Our codes and examples is availble at: https://github.com/WangTooNaive/MozartsTouch
AIGC, Multi-modal, Large Language Model, Music Generation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Sound and music computing
Figure 1. The overview of Mozart’s Touch.
1. Introduction
In recent years, the intersection of artificial intelligence (AI) and creative arts has witnessed remarkable advancements (Cao et al., 2023), leading to the emergence of novel techniques and systems capable of producing music(Schneider et al., 2023; Agostinelli et al., 2023; Copet et al., 2024), images(Rombach et al., 2022; Ramesh et al., 2021, 2022), and other forms of artistic expression(Liu et al., 2024) in a wide range of industries. As the remarkable advancements in Artificial Intelligence for Generative Composition (AIGC), there is a growing belief that it heralds a new era in AI and will have a substantial influence across the globe.
However, current music generation models, when tasked with image-to-music synthesis, encounter notable limitations. These models often struggle to accurately capture the ambiance and underlying emotions conveyed by the visual input. While they may produce music that aligns with the visual elements, the nuanced details and subtle cues present in the image are frequently lost in translation. This shortfall hampers the ability of existing systems to truly evoke the intended atmosphere and sentiment of the imagery, thereby limiting their effectiveness in multi-modal creative endeavors. It is evident that there exists a gap in the current state-of-the-art models concerning their proficiency in leveraging visual cues to inform the musical composition process.
Natural language serves as a powerful intermediary, demonstrating significant potential in bridging across different sensory modalities. Designed to interact directly with human, Large language models (LLMs) are typically comprised of a vast number of parameters and trained on extensive datasets, granting them powerful comprehension and reasoning capabilities.(Gupta and Kembhavi, 2023) Harnessing these advantages, researchers have employed LLMs to achieve semantic understanding across multiple modalities.
Despite the significant strides made in AI-driven creativity, a compelling question arises: How can we harness the formidable capabilities of LLMs to empower multi-modal tasks such as image-to-music synthesis? This inquiry serves as the focal point of our investigation, wherein we seek to elucidate the seamless integration of LLMs into the process of generating music inspired by visual contents.
In this paper, we present Mozart’s Touch, a multi-modal music generation framework that harnesses the power of Large Language Models (LLMs) and pre-trained models to generate music based on visual information. An overview of the architecture is depicted in Figure 1.
Mozart’s Touch offers multiple advantages for image-to-music generation: By leveraging the deep understanding and generalizable knowledge of Large Language Models (LLMs) to interpret visual elements accurately, it differs from previous multi-modal end-to-end music generation methods (e.g. CoDi (Tang et al., 2023) and M2UGen (Hussain et al., 2023a)). Unlike traditional approaches, it requires no training of music generation models or fine-tuning LLMs, conserving computational resources and ensuring efficiency. Moreover, Mozart’s Touch utilizes clear, interpretable prompts for greater transparency during the whole process, which improves overall framework explainability.
Our contributions are summarized as follows:
  • We introduce the Mozart’s Touch framework, an innovative integration of Large Language Models (LLMs) for multi-modal music generation. Departing from traditional end-to-end paradigms, this framework harnesses the power of LLMs to synthesize music aligned with visual inputs.
  • We offer a new perspective on leveraging LLMs for multi-modal generation tasks. Our framework showcases a novel application of LLMs in text-to-music generation , demonstrating the potential of LLMs in understanding and bridging different sensory modalities and empowering creative processes.
  • We assess Mozart’s Touch on the image- and video-to-audio dataset MUImage and MUVideo (Hussain et al., 2023b) , utilizing both objective and subjective metrics. Comparative evaluation results show that our approach outperforms existing state-of-the-art methods. This experiment demonstrates the effectiveness of our framework and its potential as a new baseline benchmark for future works in the domain.
2. Related Work
2.1. Multi-modal Large Language Model (MLLM)
Due to the prevalence of researches in Large Language Models(LLM), the combination of LLM and models in other modalities has also been a rising research hot spot, leading to the new field of MLLM. According to this survey (Yin et al., 2023) , the key applications of MLLM includes Multi-modal Instruction Tuning (M-IT), Multi-modal In-Context Learning (M-ICL), Multi-modal Chain of Thought (M-CoT), and LLM-Aided Visual Reasoning (LAVR). For Mozart’s Touch, we employ Modality Bridging technology, utilizing natural language as an intermediary medium and leveraging LLM to bridge the modality gap. VideoChat-Text (Li et al., 2023), for example, is an end-to-end chat-centric video understanding system, which uses pre-trained vision models to extract visual information such as actions and enriches the descriptions using a speech recognition model, which are all represented as textual information as a bridge.
2.2. Image Captioning
Image captioning, which is the process of generating descriptive text (captions) that accurately and relevantly capture the content of an image, is a typical multi-modal task requiring both abilities of visual understanding and natural language generation. (Stefanini et al., 2022) The field of image captioning has seen significant advancements, such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022) model. CLIP is developed by OpenAI that has revolutionized the way computers understand images and text, which efficiently learns visual concepts from natural language supervision. The main idea of CLIP is to align texts and images in the feature domain without predetermined labels for specific object categories by training on a large corpus of image-text pairs collected from the Internet. BLIP is another multi-modal framework which transfers flexibly to both vision-language understanding and generation tasks. To pre-train a unified model with both understanding and generation capabilities, they propose multi-modal mixture of encoder-decoder (MED) and achieve great performance across multiple tasks, such as image captioning.
2.3. Multi-Modal Music Generation
The advent of Transformer and diffusion models has promoted the development of music generation models. Many impressive works emerged in recent years, such as MusicLM (Agostinelli et al., 2023), MusicGen (Copet et al., 2024) , Noise2Music (Huang et al., 2023) and AudioLDM 2 (Liu et al., 2023a) . MusicLM and MusicGen both consist of autoregressive decoder to generate music. MusicLM can generate high-quality music based on descriptive text such as emotions, styles and instruments. Noise2Music and AudioLDM 2 use diffusion models to generate music based on text that transcends fine-grained semantics and can reach deeper emotions.
However, these works above all take text or audio as input to generate music, ignoring other modality information, such as image and video. Notable exceptions include the CoDi (Tang et al., 2023) and M2UGen (Hussain et al., 2023b), which allow inputs with more modalities. CoDi(Composable Diffusion) can generate output modalities in parallel from any combination of input modalities. It first use individual modality-specific diffusion models for images, videos, audio, and texts respectively to build a shared multimodal space, and then uses Latent Alignment (Deng et al., 2018) to achieve joint multi-modal generation. M2UGen is an LLM-based multi-modal music understanding and generation framework. It consists of multi-modal feature encoders, multi-model understanding adapters, bridging LLM, and generation modules to process inputs from multiple modalities such as text, images, and videos, and generate corresponding music.
3. Mozart’s Touch
Mozart’s Touch is a collaborative multi-modal AIGC framework structured into a sequential integration of three core modules: a Multi-modal Captioning Module, a LLM Understanding & Bridging Module based on LLMs and Music Generation Module. The overall architecture is illustrated in Figure 1.
3.1. Multi-modal Captioning Module
The Multi-modal Captioning Module is responsible to encode and understand users’ input, providing textual descriptions for multi-modality. This module employs state-of-the-art techniques ViT (Dosovitskiy et al., 2020) and BLIP (Li et al., 2022) model to analyze images and videos and generate descriptive captions. When users input images and videos without prompting, Our framework can also performs well to generate music that aptly complements the theme. However, in consideration of customization, we also permit users to input textual prompts to guide the music generation process.
3.1.1. Image Captioning Process
For image inputs, we leverage the capabilities of Vision Transformer (ViT) and BLIP-base modules, implemented by the clip-interrogator, to analyze and generate descriptions of the images. This process involves interpreting the visual content of an image $I$ and converting it into a image caption description $D_{\text{caption}}$.
Given an input image $I$, the framework generates a caption description $D_{\text{caption}}$ :
(1) $D_{\text{caption}}=f_{\text{BLIP}}(I)$
where $f_{\text{BLIP}}$ denotes the BLIP model to convert images into descriptive texts. The generated image caption description $D_{\text{caption}}$ serves as input for the subsequent process.
3.1.2. Video Process
For video inputs, we employ a two-step process to handle and interpret the content. Initially, Video-BLIP2-Preprocessor tool is used to sample frames from the video $V$, generating a set of frames $\{F_{i}\}$. Each frame $F_{i}$ is then processed to generate a textual description $D_{i}$ using the BLIP model, similar to the image process. This process can be formulated as:
(2) $\{D_{i}\}=\{f_{\text{BLIP}}(F_{i})\}$
where $f_{\text{BLIP}}$ denotes the BLIP model to convert frames into descriptive texts.
Subsequently, to synthesize a video caption description $D_{\text{caption}}$ of the entire video, we aggregate the frame descriptions $\{D_{i}\}$ and process them through Large Language Models (LLMs) to interpret and condense the video’s visual and thematic content into a coherent textual representation. This process can be represented as:
(3) $D_{\text{caption}}=f_{\text{LLM}}(\{D_{i}\}|Prompt_{video})$
where $f_{\text{LLM}}$ denotes the LLM to integrate and interpret the set of frame descriptions into a single video description $D_{\text{caption}}$ The prompt used in this process is shown in Table 1.
Table 1. Prompt template used to integrate the set of frame descriptions into video description.
Role Content
system You are about to process a sequence of captions, each corresponding to a distinct frame sampled from a video. Your task is to convert these captions into a cohesive, well-structured paragraph. This paragraph should describe the video in a fluid, engaging manner and follows these guidelines: avoiding semantic repetition to the greatest extent, and giving a description in less than 200 characters.
This video caption description $D_{\text{caption}}$ then serves as the input for subsequent process, similar to the image captioning process.
3.2. LLM Understanding & Bridging Module
LLM Understanding & Bridging Module plays a pivotal role in the transition from visual to auditory art forms. It is tasked with converting the image/video-descriptive caption text, generated by the Multi-modal Captioning Module, into prompts which are useful in musical generation. This conversion leverages the capabilities of Large Language Models (LLMs) to interpret the underlying mood, themes, and elements conveyed in the textual descriptions of images or videos.
Why we undertake the step of LLM-Bridge Module? This is because we contend that although multi-modal caption description have already been presented by Multi-modal Captioning Module, the problems of heterogeneous representations among different modalities still remain unsolved. For example, image captioning model (such as BLIP) intend to generate textual representations which lean more towards describing visual attributes (e.g. appearance, shape, etc.) while for music generation models (e.g. MusicGen), input descriptions that describe musical styles, moods and genres can lead to a better generation of music. From this prospective, we propose LLM Understanding & Bridging Module to align the two types of descriptions mentioned above.
To enhance the specificity and relevance of the generated music, the module also optimizes the prompts with additional constraints aimed at music generation. This includes specifying the music genre and incorporating several few-shot examples provided by MusicGen. The optimization process ensures that the final music-descriptive prompt $D_{\text{music}}$ not only reflects the mood and theme indicated by the input visuals but also adheres to the stylistic and genre-specific guidelines necessary for generating contextually appropriate music pieces. Two type of $Prompt_{bridge}$, for image and video input separately, are shown in Table 2 and 3
The process can be formulated as below. Given an visual descriptive caption $D_{\text{caption}}$, the module generates a corresponding music-descriptive prompt $D_{\text{music}}$ :
(4) $D_{\text{music}}=f_{\text{LLM}}(D_{\text{caption}}|Prompt_{bridge})$
where $f_{\text{LLM}}$ denotes the LLM to transform the descriptive texts into a coherent musical prompt that encapsulates the intended mood, themes, and potentially, the genre of the music to be generated, with the help of $Prompt_{bridge}$.
Table 2. Prompt template for image-to-music generation.
Role Content
system Convert in less than 200 characters this image caption to a very concise musical description with musical terms, so that it can be used as a prompt to generate music through AI model, strictly in English. If user provides prompt, give priority to information provided by user. You need to speculate the mood of the given image caption and add it to the music description. You also need to specify a music genre in the description such as pop, hip hop, funk, electronic, jazz, rock, metal, soul, R&B etc.
user a city with a tower and a castle in the background, a detailed matte painting, art nouveau, epic cinematic painting, kingslanding
assistant A grand orchestral arrangement with thunderous percussion, epic brass fanfares, and soaring strings, creating a cinematic atmosphere fit for a heroic battle.
user a group of people sitting on a beach next to a body of water, tourist destination, hawaii
assistant Pop dance track with catchy melodies, tropical percussion, and upbeat rhythms, perfect for the beach
Table 3. Prompt template for video-to-music generation.
Role Content
system Convert in less than 200 characters this video caption to a very concise musical description with musical terms, so that it can be used as a prompt to generate music through AI model, strictly in English. You need to speculate the mood of the given video caption and add it to the music description. You also need to specify a music genre in the description such as pop, hip hop, funk, electronic, jazz, rock, metal, soul, R&B etc.
user Two men playing cellos in a room with a piano and a grand glass window backdrop.
assistant Classical chamber music piece featuring cello duet, intricate piano accompaniment, the rich harmonies blend seamlessly in an elegant and refined setting, creating a symphonic masterpiece.
user A man with guitar in hand, captivates a large audience on stage at a concert. The crowd watches in awe as the performer delivers a stellar musical performance.
assistant Rock concert with dynamic guitar riffs, precise drumming, and powerful vocals, creating a captivating and electrifying atmosphere, uniting the audience in excitement and musical euphoria.
By invoking LLMs through API, the model is able to distinguish semantic nuances with high accuracy while ensuring its lightweight nature. This capability not only fosters streamlined processing but also facilitates seamless deployment of model services on servers with constrained computational resources.
3.3. Music Generation Module
The Music Generation Module utilizes the pre-trained model MusicGen-medium (Copet et al., 2024) to generate music pieces based on the music-descriptive prompts provided by the LLM Understanding & Bridging Module. MusicGen is designed to produce high-quality music compositions while accommodating various musical styles and preferences. By integrating MusicGen into the Mozart’s Touch framework, we ensure that the generated music aligns closely with the intended mood and theme extracted from the input visuals.
Given a music-descriptive prompt $D_{\text{music}}$, the Music Generation Module generates a music piece $M$:
(5) $M=f_{\text{MusicGen}}(D_{\text{music}})$
where $f_{\text{MusicGen}}$ represents the MusicGen model to transform the music prompt into music composition audio. It encapsulates the complex process of interpreting the prompts and translating them into musical elements such as melody, harmony, rhythm, and texture, ensuring that the generated music pieces accurately reflect the intended mood and themes conveyed by the input visuals.
4. Experiments
In this section, we assess the image-to-music and video-to-music generation capacities of Mozart’s Touch, with the discussion of two evaluation datasets MUImage and MUVideo, and the evaluation metrics utilized. The result of evaluation shows our current state-of-the-art performance in the task of multi-modal music generation.
4.1. Evaluation Dataset
To assess our framework’s performance of image-to-music generation, we utilize the MUImage dataset proposed by M2UGen (Hussain et al., 2023a). MUImage is assembled by obtaining music samples from the AudioSet (Gemmeke et al., 2017) with corresponding images, which contains 9,966 music-image pairs in total. We sampled 2,500 music-image pairs randomly from MUImage as our evaluation dataset.
For video-to-music generation task, we utilize the MUVideo dataset, which is also proposed by M2UGen. We adopted a construction method similar to that of the image-to-music generation task, yielding a corpus of 2,500 music-video pairs for evaluating video-to-music generation task.
4.2. Evaluation metrics
For both tasks, we utilize the Frechet Audio Distance (FAD)(Kilgour et al., 2019), Kullback-Leibler divergence (KL) and ImageBind Rank (IB Rank)(Girdhar et al., 2023) as the evaluation metrics. FAD is a reference-free evaluation metric for music enhancement algorithms. A low score of FAD indicates a high quality of generated music. KL scores measure the labels between the original and the generated music. When the KL score is low, the generated audios are expected to share similar distributions with the reference music. For these two metrics, we utilize the official implementation in PyTorch, where FAD score is supported by the VGGish model.
IB Rank(Girdhar et al., 2023) is introduced by M2UGen, to assess the alignment between the image/video modality and the generated music. Firstly, we use the Image-Bind model to obtain embeddings for the images/videos and the generated music, then calculate their cosine similarity scores and give them a score based on their ranking. For IB Rank, High score represents a relatively high ranking among the baselines.
4.3. Baselines and Details
For both tasks, we compare Mozart’s Touch with two baselines: CoDi(Tang et al., 2023) and M2UGen(Hussain et al., 2023a). We use open-source CoDi model and M2UGen checkpoint files to run inference. Our framework runs on one NVIDIA RTX 3090 24GB GPU, and two baselines run on one NVIDIA V100 32GB GPU to load the whole models.
4.4. Performance Comparison
Table 4 presents the performance of our framework, Mozart’s Touch, and two baseline models in image-to-music generation. The results highlight significant improvements in both the quality and relevance of the music generated by our framework. Moreover, Mozart’s Touch surpasses prior state-of-the-art models despite its simpler architecture.
Table 5 shows the results of video-to-music generation. For this task, we observed that Mozart’s Touch still outperforms other models, indicating that our two-step captioning strategy is also highly effective.
Table 4. Objective comparison of models for image-to-music generation. The best results are made bold.
Model $FAD_{vgg}$ KL↓ IM Rank↑
M2UGen 9.166 1.870 0.556
CoDi 6.674 1.821 0.525
Mozart’s Touch 4.625 1.169 0.753
Table 5. Objective comparison of models for video-to-music generation. The best results are made bold.
Model $FAD_{vgg}$ KL↓ IM Rank↑
M2UGen 9.047 1.878 0.552
CoDi 5.055 1.195 0.494
Mozart’s Touch 4.339 1.048 0.787
4.5. Subjective Evaluation
Although we achieve exceptional performance in the objective evaluation, we also believe that quantitative evaluation method has great limitations for music generation tasks. The metrics above can effectively measure the quality and relevance of the generated music, but fall short in the understanding of creativity and human feelings, as supported by previous research (Liu et al., 2023b).
Following previous similar works (Liu et al., 2023b; Kreuk et al., 2023), the generated samples are rated based on i) overall quality (OVL); and ii) relevance to the input image (REL). Both OVL and REL metrics have a Likert scale (Likert, 1932) between one and five, where a larger number indicates better performance.
In this case, We conduct the subjective evaluation involving 125 participants, taking image-to-music generation as example. Totally 75 questions are created for the subjective evaluation, which are randomly sampled from our evaluation dataset. Each question contains a video with the input image as the visual part and generated (or ground truth) music as the audio. 20 audios are sampled from ground truth, 20 from M2UGen, 20 from Mozart’s Touch, and 15 from CoDi. Each questionnaire comprises ten randomly selected questions. Upon subsequent validation by our team, all 75 questions are covered by the total 125 questionnaires.
Table 6. Subjective comparison of models for image-to-music generation. The best results are made bold.
Model OVL↑ REL↑
CoDi 2.95 3.24
M2UGen 3.77 3.02
Mozart’s Touch 3.74 3.76
Ground Truth 3.88 4.08
The subjective evaluation result is presented in Table 6. While our method slightly underperforms in terms of the metrics for overall quality (OVL) when compared to M2UGen, the result shows that there is a notable enhancement in the metric of relevance (REL) to input image, which is consistent with our target to generate corresponding music that aligns the image well.
4.6. Ablation Studies
Table 7. Ablation study on image-to-music generation task. The best results are made bold.
Model $FAD_{vgg}$ KL↓ IM Rank↑
Mozart’s Touch 4.625  1.170 0.757
w/o LUBM 3.741 1.121 0.743
To demonstrate the effectiveness of LLM bridging modality, we conducted a further ablation experiment, comparing the performance of the original system with and without (w/o) the LLM Understanding & Bridging Module (LUBM) in the task of iamge-to-music generation.
As indicated in the table 7, the framework without LUBM achieves higher scores in the FAD and KL metrics, the two metrics measure the similarity between ground truth and generated audios, rather than the similarity between different modalities. On the other side, the framework with LUBM performs better in IB Rank metric. This metric utilizes the ImageBind model to encode multi-modal information uniformly, thereby evaluating the similarity between input modality information and generated audio, aligning more closely with the objectives of evaluating multi-modal music generation. Therefore, we believe that there is no clear superiority or inferiority between the Mozart’s Touch framework with and without LUBM. This once again emphasizes that quantitative evaluation may not always be the best approach for assessing the multi-modal music generation tasks.
4.7. Case Study
In this part, we conduct a case study to analyze how our LLM Understanding & Bridging Module (LUBM) mitigates the problem of heterogeneous representations among different modalities. By showcasing some representative comparative examples in Figure 2, We demonstrate that the absence of the LUBM does indeed have adverse effects on the generation results.
The first example illustrates a portrait of Bach. Some keywords in the original image description disturb the generation of corresponding music, as they focus on the attributes of image instead of that of music. The second example illustrates an anime girl from a visual novel game Atri: My Dear Moments. This example shows that insufficiency of music attributions may also mislead the generation of music in a quite different way.
Figure 2. Case Study
5. Conclusion
This paper introduces Mozart’s Touch, a lightweight multi-modal music generation framework that seamlessly integrates LLMs with pre-trained models together. Experiments and researches demonstrate the framework’s capability to perform multi-modality understanding and captioning, multi-modal representations bridging, and music generation, resulting in highly aligned music based on the corresponding multi-modal inputs.
Our future work will aim to refine our prompting strategy for enhanced alignment with multi-modal inputs, conduct further evaluation experiments on our LLM Understanding & Bridging Module, and integrate recent advancements actively into our framework. Meanwhile, we will maintain its lightweight characteristic of our framework, ensuring its user-friendliness and expanding its accessibility to a broader scenario.

### v3

@@
-Mozart’s Touch: A Lightweight Multi-modal Music Generation Framework Based on Pre-Trained Large Models
-Abstract.
-In recent years, AI-Generated Content (AIGC) has witnessed rapid advancements, facilitating the generation of music, images, and other forms of artistic expression across various industries.
-However, researches on general multi-modal music generation model remain scarce.
-To fill this gap, we propose a multi-modal music generation framework Mozart’s Touch.
-It could generate aligned music with the cross-modality inputs, such as images, videos and text.
-Mozart’s Touch is composed of three main components: Multi-modal Captioning Module, Large Language Model (LLM) Understanding & Bridging Module, and Music Generation Module.
-Unlike traditional approaches, Mozart’s Touch requires no training or fine-tuning pre-trained models, offering efficiency and transparency through clear, interpretable prompts.
-We also introduce ”LLM-Bridge” method to resolve the heterogeneous representation problems between descriptive texts of different modalities.
-We conduct a series of objective and subjective evaluations on the proposed model, and results indicate that our model surpasses the performance of current state-of-the-art models.
-Our codes and examples is availble at: https://github.com/WangTooNaive/MozartsTouch
-AIGC, Multi-modal, Large Language Model, Music Generation
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Sound and music computing
-Figure 1. The overview of Mozart’s Touch.
-1. Introduction
-In recent years, the intersection of artificial intelligence (AI) and creative arts has witnessed remarkable advancements (Cao et al., 2023), leading to the emergence of novel techniques and systems capable of producing music(Schneider et al., 2023; Agostinelli et al., 2023; Copet et al., 2024), images(Rombach et al., 2022; Ramesh et al., 2021, 2022), and other forms of artistic expression(Liu et al., 2024) in a wide range of industries.
-As the remarkable advancements in Artificial Intelligence for Generative Composition (AIGC), there is a growing belief that it heralds a new era in AI and will have a substantial influence across the globe.
+Further author information:
+Abstract
+In recent years, AI-Generated Content (AIGC) has witnessed rapid advancements, facilitating the creation of music, images, and other artistic forms across a wide range of industries.
+However, current models for image- and video-to-music synthesis struggle to capture the nuanced emotions and atmosphere conveyed by visual content.
+To fill this gap, we propose Mozart’s Touch, a multi-modal music generation framework capable of generating music aligned with cross-modal inputs such as images, videos, and text.
+The framework consists of three key components: Multi-modal Captioning Module, Large Language Model (LLM) Understanding & Bridging Module, and Music Generation Module.
+Unlike traditional end-to-end methods, Mozart’s Touch uses LLMs to accurately interpret visual elements without requiring the training or fine-tuning of music generation models, providing efficiency and transparency through clear, interpretable prompts.
+We also introduce the ”LLM-Bridge” method to resolve the heterogeneous representation challenges between descriptive texts from different modalities.
+Through a series of objective and subjective evaluations, we demonstrate that Mozart’s Touch outperforms current state-of-the-art models.
+Our code and examples are available at https://github.com/TiffanyBlews/MozartsTouch.
+keywords:
+AIGC, Multi-modal, Neural Network, Large Language Model, Music Generation
+Figure 1: The overview of Mozart’s Touch.
+1 Introduction
+In recent years, the intersection of artificial intelligence (AI) and creative arts has led to significant advancements [1], leading to the emergence of novel techniques and systems capable of producing music[2, 3, 4], images[5, 6, 7], and other artistic forms[8] .
+These developments, particularly in Artificial Intelligence for Generative Composition (AIGC), are widely believed to mark the beginning of a new era in AI, with far-reaching global implications.
 However, current music generation models, when tasked with image-to-music synthesis, encounter notable limitations.
-These models often struggle to accurately capture the ambiance and underlying emotions conveyed by the visual input.
-While they may produce music that aligns with the visual elements, the nuanced details and subtle cues present in the image are frequently lost in translation.
-This shortfall hampers the ability of existing systems to truly evoke the intended atmosphere and sentiment of the imagery, thereby limiting their effectiveness in multi-modal creative endeavors.
-It is evident that there exists a gap in the current state-of-the-art models concerning their proficiency in leveraging visual cues to inform the musical composition process.
-Natural language serves as a powerful intermediary, demonstrating significant potential in bridging across different sensory modalities.
-Designed to interact directly with human, Large language models (LLMs) are typically comprised of a vast number of parameters and trained on extensive datasets, granting them powerful comprehension and reasoning capabilities.(Gupta and Kembhavi, 2023) Harnessing these advantages, researchers have employed LLMs to achieve semantic understanding across multiple modalities.
-Despite the significant strides made in AI-driven creativity, a compelling question arises: How can we harness the formidable capabilities of LLMs to empower multi-modal tasks such as image-to-music synthesis?
-This inquiry serves as the focal point of our investigation, wherein we seek to elucidate the seamless integration of LLMs into the process of generating music inspired by visual contents.
-In this paper, we present Mozart’s Touch, a multi-modal music generation framework that harnesses the power of Large Language Models (LLMs) and pre-trained models to generate music based on visual information.
-An overview of the architecture is depicted in Figure 1.
-Mozart’s Touch offers multiple advantages for image-to-music generation: By leveraging the deep understanding and generalizable knowledge of Large Language Models (LLMs) to interpret visual elements accurately, it differs from previous multi-modal end-to-end music generation methods (e.g.
-CoDi (Tang et al., 2023) and M2UGen (Hussain et al., 2023a)).
-Unlike traditional approaches, it requires no training of music generation models or fine-tuning LLMs, conserving computational resources and ensuring efficiency.
-Moreover, Mozart’s Touch utilizes clear, interpretable prompts for greater transparency during the whole process, which improves overall framework explainability.
+These models often struggle to accurately capture the ambiance and underlying emotions conveyed by visual input.
+While the generated music may align with the visual elements, the nuanced details and subtle cues embedded in the image are often lost in translation, preventing the music from fully reflecting the intended atmosphere and sentiment.
+This limitation highlights a significant gap in current state-of-the-art models concerning their proficiency in leveraging visual cues to inform the musical composition process.
+Natural language has emerged as a powerful bridge between different sensory modalities, enabling deeper semantic understanding.
+Designed to interact directly with human, Large Language Models (LLMs) are typically comprised of vast numbers of parameters and trained on extensive datasets, granting them powerful comprehension and reasoning capabilities.[9] Harnessing these advantages, researchers have employed LLMs to achieve semantic understanding across multiple modalities.
+Despite the significant strides made in AI-driven creativity, a critical question remains: How can we leverage the full potential of LLMs to empower multi-modal tasks such as image-to-music synthesis?
+This paper aims to explore this question by presenting a novel framework that integrates LLMs into the music generation process, inspired by visual content.
+We introduce Mozart’s Touch, a multi-modal music generation framework that utilizes LLMs and pre-trained models to generate music based on visual information.
+An overview of the architecture is illustrated in Fig.
+1.
+Mozart’s Touch offers several key advantages for image- and video-to-music generation: Unlike previous end-to-end multi-modal methods (e.g., CoDi [10] and M2UGen [11]), our approach leverages the deep understanding and generalizable knowledge of LLMs to interpret visual elements accurately, without the need for training new music generation models or fine-tuning LLMs.
+This not only conserves computational resources but also ensures greater efficiency.
+Furthermore, Mozart’s Touch employs clear and interpretable prompts, enhancing the transparency and explainability of the process.
 Our contributions are summarized as follows:
-  • We introduce the Mozart’s Touch framework, an innovative integration of Large Language Models (LLMs) for multi-modal music generation.
-Departing from traditional end-to-end paradigms, this framework harnesses the power of LLMs to synthesize music aligned with visual inputs.
-  • We offer a new perspective on leveraging LLMs for multi-modal generation tasks.
-Our framework showcases a novel application of LLMs in text-to-music generation , demonstrating the potential of LLMs in understanding and bridging different sensory modalities and empowering creative processes.
-  • We assess Mozart’s Touch on the image- and video-to-audio dataset MUImage and MUVideo (Hussain et al., 2023b) , utilizing both objective and subjective metrics.
-Comparative evaluation results show that our approach outperforms existing state-of-the-art methods.
-This experiment demonstrates the effectiveness of our framework and its potential as a new baseline benchmark for future works in the domain.
-2. Related Work
-2.1. Multi-modal Large Language Model (MLLM)
+  • We present Mozart’s Touch, a novel framework that integrates LLMs for multi-modal music generation.
+By departing from traditional end-to-end paradigms, this framework uses LLMs to generate music that aligns with visual inputs.
+  • We evaluate Mozart’s Touch using the image- and video-to-audio datasets MUImage and MUVideo [12], applying both objective and subjective metrics.
+The results demonstrate that our framework outperforms existing state-of-the-art methods, establishing it as a promising new benchmark in the field.
+  • We offer a fresh perspective on leveraging LLMs for multi-modal tasks, showcasing their potential in bridging different sensory modalities to empower the creative process, particularly in text-to-music generation.
+2 Related Work
+2.1 Multi-modal Large Language Model (MLLM)
 Due to the prevalence of researches in Large Language Models(LLM), the combination of LLM and models in other modalities has also been a rising research hot spot, leading to the new field of MLLM.
-According to this survey (Yin et al., 2023) , the key applications of MLLM includes Multi-modal Instruction Tuning (M-IT), Multi-modal In-Context Learning (M-ICL), Multi-modal Chain of Thought (M-CoT), and LLM-Aided Visual Reasoning (LAVR).
+According to this survey [13] , the key applications of MLLM includes Multi-modal Instruction Tuning (M-IT), Multi-modal In-Context Learning (M-ICL), Multi-modal Chain of Thought (M-CoT), and LLM-Aided Visual Reasoning (LAVR).
 For Mozart’s Touch, we employ Modality Bridging technology, utilizing natural language as an intermediary medium and leveraging LLM to bridge the modality gap.
-VideoChat-Text (Li et al., 2023), for example, is an end-to-end chat-centric video understanding system, which uses pre-trained vision models to extract visual information such as actions and enriches the descriptions using a speech recognition model, which are all represented as textual information as a bridge.
-2.2. Image Captioning
+VideoChat-Text [14], for example, is an end-to-end chat-centric video understanding system, which uses pre-trained vision models to extract visual information such as actions and enriches the descriptions using a speech recognition model, which are all represented as textual information as a bridge.
+2.2 Image Captioning
 Image captioning, which is the process of generating descriptive text (captions) that accurately and relevantly capture the content of an image, is a typical multi-modal task requiring both abilities of visual understanding and natural language generation.
-(Stefanini et al., 2022) The field of image captioning has seen significant advancements, such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022) model.
+[15] The field of image captioning has seen significant advancements, such as CLIP [16] and BLIP [17] model.
 CLIP is developed by OpenAI that has revolutionized the way computers understand images and text, which efficiently learns visual concepts from natural language supervision.
@@
 To pre-train a unified model with both understanding and generation capabilities, they propose multi-modal mixture of encoder-decoder (MED) and achieve great performance across multiple tasks, such as image captioning.
-2.3. Multi-Modal Music Generation
+2.3 Multi-Modal Music Generation
 The advent of Transformer and diffusion models has promoted the development of music generation models.
-Many impressive works emerged in recent years, such as MusicLM (Agostinelli et al., 2023), MusicGen (Copet et al., 2024) , Noise2Music (Huang et al., 2023) and AudioLDM 2 (Liu et al., 2023a) .
+Many impressive works emerged in recent years, such as MusicLM [3], MusicGen [4] , Noise2Music [18] and AudioLDM 2 [19] .
 MusicLM and MusicGen both consist of autoregressive decoder to generate music.
@@
 However, these works above all take text or audio as input to generate music, ignoring other modality information, such as image and video.
-Notable exceptions include the CoDi (Tang et al., 2023) and M2UGen (Hussain et al., 2023b), which allow inputs with more modalities.
+Notable exceptions include the CoDi [10] and M2UGen [12], which allow inputs with more modalities.
 CoDi(Composable Diffusion) can generate output modalities in parallel from any combination of input modalities.
-It first use individual modality-specific diffusion models for images, videos, audio, and texts respectively to build a shared multimodal space, and then uses Latent Alignment (Deng et al., 2018) to achieve joint multi-modal generation.
+It first use individual modality-specific diffusion models for images, videos, audio, and texts respectively to build a shared multimodal space, and then uses Latent Alignment [20] to achieve joint multi-modal generation.
 M2UGen is an LLM-based multi-modal music understanding and generation framework.
 It consists of multi-modal feature encoders, multi-model understanding adapters, bridging LLM, and generation modules to process inputs from multiple modalities such as text, images, and videos, and generate corresponding music.
-3. Mozart’s Touch
+3 Mozart’s Touch
 Mozart’s Touch is a collaborative multi-modal AIGC framework structured into a sequential integration of three core modules: a Multi-modal Captioning Module, a LLM Understanding & Bridging Module based on LLMs and Music Generation Module.
-The overall architecture is illustrated in Figure 1.
-3.1. Multi-modal Captioning Module
-The Multi-modal Captioning Module is responsible to encode and understand users’ input, providing textual descriptions for multi-modality.
-This module employs state-of-the-art techniques ViT (Dosovitskiy et al., 2020) and BLIP (Li et al., 2022) model to analyze images and videos and generate descriptive captions.
-When users input images and videos without prompting, Our framework can also performs well to generate music that aptly complements the theme.
-However, in consideration of customization, we also permit users to input textual prompts to guide the music generation process.
-3.1.1. Image Captioning Process
-For image inputs, we leverage the capabilities of Vision Transformer (ViT) and BLIP-base modules, implemented by the clip-interrogator, to analyze and generate descriptions of the images.
-This process involves interpreting the visual content of an image $I$ and converting it into a image caption description $D_{\text{caption}}$.
-Given an input image $I$, the framework generates a caption description $D_{\text{caption}}$ :
-(1) $D_{\text{caption}}=f_{\text{BLIP}}(I)$
-where $f_{\text{BLIP}}$ denotes the BLIP model to convert images into descriptive texts. The generated image caption description $D_{\text{caption}}$ serves as input for the subsequent process.
-3.1.2. Video Process
-For video inputs, we employ a two-step process to handle and interpret the content.
-Initially, Video-BLIP2-Preprocessor tool is used to sample frames from the video $V$, generating a set of frames $\{F_{i}\}$.
+The overall architecture is illustrated in Fig.
+1.
+Specifically, the process is shown in Algorithm 1 in pseudo-code.
+Initially, the Multi-modal Captioning Module processes the input visuals, which may be either images or videos, using the BLIP model [17].
+If the input is an image, a single caption is generated.
+If the input is a video, the algorithm samples frames from the video and generates individual captions for each frame.
+These frame-level captions are then aggregated into a comprehensive description using an LLM, forming a video-level caption.
+In the subsequent LLM Understanding & Bridging Module, the captions, whether image-based or video-based, are fed into a large language model (LLM).
+This model transforms the captions into a music-descriptive prompt tailored for the respective visual input.
+For images, the transformation is guided by an image-specific bridging prompt, while for videos, the transformation uses a video-specific prompt.
+The final stage, the Music Generation Module, uses the transformed music-descriptive prompt to generate a piece of music with the pre-trained MusicGen model [4].
+The output is the generated music, which is returned as the final result of the pipeline.
+3.1 Multi-modal Captioning Module
+The Multi-modal Captioning Module is responsible for processing and understanding user input, providing textual descriptions from multi-modal data.
+This module employs state-of-the-art techniques ViT [21] and BLIP [17] model to analyze images and videos and generate descriptive captions.
+When users input images or videos without specific prompts, Our framework still generates music that appropriately complements the underlying theme.
+However, in consideration of customization, users may also input textual prompts to guide the music generation process.
+3.1.1 Image and Video Captioning Process
+For both image and video inputs, we leverage the capabilities of Vision Transformer (ViT) and BLIP model, implemented by the clip-interrogator, to analyze and generate descriptions of the images.
+When an image $I$ is provided, the BLIP model directly generates a caption $C_{I}$ by processing the image:
+$C_{I}=\textsc{GenerateCaption}(I,\text{BLIP})$ (1)
+Algorithm 1 Multi-modal to Music Generation Framework
+  Input: Visual inputs $V_{input}$ (image or video), optional textual prompts $T$
+  Output: Generated music $M$
+  // Module 1: Multi-modal Captioning
+  if $V_{input}$ is an image $I$ then
+     $C_{I}\leftarrow\textsc{GenerateCaption}(I,\text{BLIP})$
+  else if $V_{input}$ is a video $V$ then
+     $\{F_{i}\}\leftarrow\textsc{SampleFrames}(V,\text{VideoProcessor})$
+     $C_{F}\leftarrow\emptyset$
+     for all $F_{i}\in\{F_{i}\}$ do
+        $C_{F}\leftarrow C_{F}\cup\textsc{GenerateCaption}(F_{i},\text{BLIP})$
+     end for
+     $C_{V}\leftarrow\textsc{AggregateCaptions}(C_{F},\text{LLM},\text{Prompt}_{video})$
+  end if
+  // Module 2: LLM Understanding & Bridging
+  if $V_{input}$ is an image $I$ then
+     $P_{M}\leftarrow\textsc{TransformCaption}(C_{I},T,\text{LLM},\text{Prompt}_{bridge}^{image})$
+  else if $V_{input}$ is a video $V$ then
+     $P_{M}\leftarrow\textsc{TransformCaption}(C_{V},T,\text{LLM},\text{Prompt}_{bridge}^{video})$
+  end if
+  // Module 3: Music Generation
+  $M\leftarrow\textsc{GenerateMusic}(P_{M},\text{MusicGen})$
+  Return: $M$
+For video inputs, the process involves two steps.
+First, the video $V$ is sampled into frames $\{F_{i}\}$ using the VideoProcessor tool.
 Each frame $F_{i}$ is then processed to generate a textual description $D_{i}$ using the BLIP model, similar to the image process.
-This process can be formulated as:
-(2) $\{D_{i}\}=\{f_{\text{BLIP}}(F_{i})\}$
-where $f_{\text{BLIP}}$ denotes the BLIP model to convert frames into descriptive texts.
-Subsequently, to synthesize a video caption description $D_{\text{caption}}$ of the entire video, we aggregate the frame descriptions $\{D_{i}\}$ and process them through Large Language Models (LLMs) to interpret and condense the video’s visual and thematic content into a coherent textual representation.
-This process can be represented as:
-(3) $D_{\text{caption}}=f_{\text{LLM}}(\{D_{i}\}|Prompt_{video})$
-where $f_{\text{LLM}}$ denotes the LLM to integrate and interpret the set of frame descriptions into a single video description $D_{\text{caption}}$ The prompt used in this process is shown in Table 1.
-Table 1. Prompt template used to integrate the set of frame descriptions into video description.
+$\{C_{F_{i}}\}=\{\textsc{GenerateCaption}(F_{i},\text{BLIP})\}$ (2)
+Next, these frame-level captions $\{C_{F_{i}}\}$ are aggregated into a single video-level caption $C_{V}$ using a Large Language Model (LLM) that interprets and synthesizes the individual frame descriptions into a cohesive narrative about the video:
+$C_{V}=\textsc{AggregateCaptions}(\{C_{F_{i}}\},\text{LLM},\text{Prompt}_{\text{video}})$ (3)
+The prompt used in this process is shown in Tab. 4 in Appendix A. This unified caption $C_{V}$ serves as the input for the subsequent stages of the pipeline, just as the image caption $C_{I}$ does.
+3.2 LLM Understanding & Bridging Module
+The LLM Understanding & Bridging Module plays a crucial role in transforming visual descriptions into music-descriptive prompts.
+It is tasked to convert the captions generated by the Multi-modal Captioning Module into prompts suitable for music generation.
+This conversion leverages Large Language Models (LLMs) to interpret the underlying mood, themes, and elements conveyed in the textual descriptions of images or videos.
+Why LLM-Bridge is Necessary?
+While the Multi-modal Captioning Module produces textual descriptions of images or videos, these descriptions are typically focused on visual elements, which may not fully align with the musical attributes required for generating music.
+The LLM Understanding & Bridging Module addresses this by interpreting and converting these visual descriptions into music-relevant prompts.
+For example, image captions often describe appearance or shape, whereas music generation requires descriptions of mood, style, and genre.
+The LLM-Bridge Module ensures that these heterogeneous representations are aligned for music generation.
+To optimize the music-descriptive prompts, the module applies additional constraints that specify music genres and other stylistic attributes.
+The goal is to ensure that the generated prompt reflects the visual input’s mood and themes while adhering to musical styles and genre-specific guidelines, improving the relevance and quality of the resulting music.
+The process can be formulated as follows: Given an input caption $C_{V}$ (for video) or $C_{I}$ (for image), with user input textual prompt $T$ as an optional, the LLM Understanding & Bridging Module generates a music-descriptive prompt $P_{M}$:
+$P_{M}=\textsc{TransformCaption}(C_{X},T,\text{LLM},\text{Prompt}_{\text{bridge}}^{X})$ (4)
+where $C_{X}$ is the caption $C_{V}$ or $C_{I}$, and $\text{Prompt}_{\text{bridge}}^{X}$ refers to the respective bridging prompt for video or image input.
+The LLM then generates a music-descriptive prompt $P_{M}$ that encapsulates the intended mood, themes, and musical attributes such as music genre.
+$\text{Prompt}_{\text{bridge}}^{X}$ for image and video input separately, are shown in Tab. 5 and 6 in Appendix A.
+3.3 Music Generation Module
+The Music Generation Module generates the final music piece using the pre-trained MusicGen model [4], based on the music-descriptive prompt $P_{M}$ generated by the LLM Understanding & Bridging Module.
+MusicGen is designed to produce high-quality compositions across various musical styles, ensuring that the generated music aligns with the intended mood and themes extracted from the input visuals.
+Given a music-descriptive prompt $P_{M}$, the Music Generation Module generates a music piece $M$:
+$M=\textsc{GenerateMusic}(P_{M},\text{MusicGen})$ (5)
+Here, the MusicGen model interprets the prompt and generates a composition that reflects the desired mood, style, and thematic elements.
+The final music piece $M$ is the output of the pipeline, effectively completing the transformation from visual input to music generation.
+4 Experiments
+In this section, we evaluate the image-to-music and video-to-music generation capabilities of Mozart’s Touch, with the discussion of two evaluation datasets (MUImage and MUVideo), the metrics used for assessment, and the results, which demonstrate the state-of-the-art performance of our framework in multi-modal music generation.
+4.1 Evaluation Dataset
+To assess the image-to-music generation capabilities of our framework, we use the MUImage dataset introduced by M2UGen [11].
+This dataset consists of 9,966 music-image pairs, derived from the AudioSet [22].
+For evaluation, we randomly sampled 2,500 music-image pairs from MUImage.
+For video-to-music generation task, we utilize the MUVideo dataset, also introduced by M2UGen.
+We adopted a construction method similar to that of the image-to-music generation task, yielding a corpus of 2,500 music-video pairs for evaluating video-to-music generation task.
+4.2 Evaluation metrics
+For both tasks, we utilize the Frechet Audio Distance (FAD)[23], Kullback-Leibler divergence (KL) and ImageBind Rank (IB Rank)[24] as the evaluation metrics.
+FAD is a reference-free evaluation metric for music enhancement algorithms.
+A low score of FAD indicates a high quality of generated music.
+KL scores measure the labels between the original and the generated music.
+When the KL score is low, the generated audios are expected to share similar distributions with the reference music.
+For these two metrics, we utilize the official implementation in PyTorch, where FAD score is supported by the VGGish model.
+IB Rank[24] is introduced by M2UGen, to assess the alignment between the image/video modality and the generated music.
+Firstly, we use the Image-Bind model to obtain embeddings for the images/videos and the generated music, then calculate their cosine similarity scores and give them a score based on their ranking.
+For IB Rank, High score represents a relatively high ranking among the baselines.
+4.3 Baselines and Experimental Setup
+We compare Mozart’s Touch against two baseline models: CoDi [10] and M2UGen [11].
+Both baselines were evaluated using their open-source implementations and pre-trained checkpoints.
+Our framework runs on one NVIDIA RTX 3090 24GB GPU, and two baselines run on one NVIDIA V100 32GB GPU to load the whole models.
+4.4 Performance Comparison
+Table 1(a) summarizes the image-to-music generation results, showing that Mozart’s Touch achieves significant improvements in both music quality and alignment with visual inputs.
+Despite its simpler architecture, our framework consistently outperforms prior state-of-the-art models.
+Table 1(b) presents the video-to-music generation results, where Mozart’s Touch also demonstrates superior performance.
+These findings validate the effectiveness of our two-step captioning strategy, highlighting its adaptability across different modalities.
+Table 1: Objective comparison of models for image-to-music and video-to-music generation. The best results are made bold.
+Model $FAD_{vgg}$ KL↓ IM Rank↑
+M2UGen 9.166 1.870 0.556
+CoDi 6.674 1.821 0.525
+Mozart’s Touch 4.625 1.169 0.753
+(a) Image-to-Music
+Model $FAD_{vgg}$ KL↓ IM Rank↑
+M2UGen 9.047 1.878 0.552
+CoDi 5.055 1.195 0.494
+Mozart’s Touch 4.339 1.048 0.787
+(b) Video-to-Music
+4.5 Subjective Evaluation
+Although we achieve exceptional performance in the objective evaluation, we also believe that quantitative evaluation method has great limitations for music generation tasks.
+The metrics above can effectively measure the quality and relevance of the generated music, but fall short in capturing creativity and the emotional resonance with human listeners, as highlighted in prior research [25].
+To address this, we conducted a subjective evaluation following methodologies from similar studies [25, 26].
+The generated samples are rated based on i) overall quality (OVL); and ii) relevance to the input image (REL).
+Both OVL and REL metrics have a Likert scale [27] ranging from 1 to 5, where higher scores indicate better performance.
+The subjective evaluation involved 125 participants, taking image-to-music generation as example.
+A total of 75 questions are created for the subjective evaluation, which are randomly sampled from our evaluation dataset.
+Each question contains a video with the input image as the visual part and generated (or ground truth) music as the audio.
+20 audios are sampled from ground truth, 20 from M2UGen, 20 from Mozart’s Touch, and 15 from CoDi.
+Each questionnaire comprises ten randomly selected questions.
+Upon subsequent validation by our team, all 75 questions are covered by the total 125 questionnaires.
+Table 2: Subjective comparison of models for image-to-music generation. The best results are made bold.
+Model OVL↑ REL↑
+CoDi 2.95 3.24
+M2UGen 3.77 3.02
+Mozart’s Touch 3.74 3.76
+Ground Truth 3.88 4.08
+The results of the subjective evaluation are presented in Tab.
+2.
+While Mozart’s Touch shows a slight underperformance in the Overall Quality (OVL) metric compared to M2UGen, it demonstrates a significant improvement in the Relevance (REL) metric, aligning more effectively with the input image, which is consistent with our primary objective of generating music that closely reflects the visual content, underscoring the strength of our framework in achieving image-to-music alignment.
+4.6 Ablation Studies
+To evaluate the effectiveness of the LLM Understanding & Bridging Module (LUBM), we conducted an ablation study comparing the performance of the complete Mozart’s Touch framework with and without LUBM in the task of image-to-music generation.
+As shown in Tab.
+3, the framework without LUBM achieves higher scores in the FAD and KL metrics.
+These metrics measure the similarity between the ground truth and the generated audio, focusing on audio quality rather than cross-modal alignment.
+Conversely, the framework with LUBM performs better in the IB Rank metric, which evaluates the alignment between visual inputs and generated audio using the ImageBind model.
+By encoding multi-modal information uniformly, IB Rank provides a more relevant assessment of cross-modal alignment, aligning with the objectives of multi-modal music generation.
+Table 3: Ablation study on image-to-music generation task. The best results are made bold.
+Model $FAD_{vgg}$ KL↓ IM Rank↑
+Mozart’s Touch 4.625  1.170 0.757
+w/o LUBM 3.741 1.121 0.743
+These results indicate that while the absence of LUBM may lead to higher scores in metrics centered on intra-modality similarity, the inclusion of LUBM enhances alignment across modalities.
+Thus, neither configuration demonstrates absolute superiority or inferiority.
+This further highlights the limitations of quantitative evaluation methods in fully capturing the complexities of multi-modal music generation tasks.
+4.7 Case Study
+To further analyze the impact of LUBM on bridging heterogeneous representations across modalities, we conducted a case study, presenting representative examples in Fig.
+2  [^1] Multimedia demonstration page at https://tiffanyblews.github.io/MozartsTouch-demo/.
+These examples illustrate how the absence of LUBM affects the quality and relevance of the generated music.
+Figure 2: Case Study
+In the first example, a portrait of Bach is used as input.
+Without LUBM, certain keywords in the image description inadvertently dominate the generation process, focusing on visual attributes instead of musical characteristics.
+This results in music that is less aligned with the essence of the visual input.
+The second example features an anime character from the visual novel game Atri: My Dear Moments.
+Without LUBM, the lack of explicit musical attributes in the image leads to a divergence in the generated music, resulting in outputs that deviate significantly from the intended alignment with the visual theme.
+These examples demonstrate how LUBM mitigates the challenges of heterogeneous representations by enabling a deeper understanding of input modalities, ensuring the generated music is both relevant and cohesive with the visual content.
+5 Conclusion
+In this paper, we introduced Mozart’s Touch, a lightweight and efficient multi-modal music generation framework that integrates Large Language Models (LLMs) with pre-trained models.
+Extensive experiments demonstrated its capability for multi-modal understanding, bridging visual and auditory modalities, and generating music highly aligned with input data.
+Future work will focus on refining the prompting strategy, conducting deeper evaluations of the LLM Understanding & Bridging Module, and integrating recent advancements.
+By maintaining its lightweight design, we aim to expand its accessibility and applicability across diverse scenarios.
+Appendix A Prompt Template
+Table 4: Prompt template used to integrate the set of frame descriptions into video description.
 Role Content
@@
 This paragraph should describe the video in a fluid, engaging manner and follows these guidelines: avoiding semantic repetition to the greatest extent, and giving a description in less than 200 characters.
-This video caption description $D_{\text{caption}}$ then serves as the input for subsequent process, similar to the image captioning process.
-3.2. LLM Understanding & Bridging Module
-LLM Understanding & Bridging Module plays a pivotal role in the transition from visual to auditory art forms.
-It is tasked with converting the image/video-descriptive caption text, generated by the Multi-modal Captioning Module, into prompts which are useful in musical generation.
-This conversion leverages the capabilities of Large Language Models (LLMs) to interpret the underlying mood, themes, and elements conveyed in the textual descriptions of images or videos.
-Why we undertake the step of LLM-Bridge Module?
-This is because we contend that although multi-modal caption description have already been presented by Multi-modal Captioning Module, the problems of heterogeneous representations among different modalities still remain unsolved.
-For example, image captioning model (such as BLIP) intend to generate textual representations which lean more towards describing visual attributes (e.g.
-appearance, shape, etc.) while for music generation models (e.g.
-MusicGen), input descriptions that describe musical styles, moods and genres can lead to a better generation of music.
-From this prospective, we propose LLM Understanding & Bridging Module to align the two types of descriptions mentioned above.
-To enhance the specificity and relevance of the generated music, the module also optimizes the prompts with additional constraints aimed at music generation.
-This includes specifying the music genre and incorporating several few-shot examples provided by MusicGen.
-The optimization process ensures that the final music-descriptive prompt $D_{\text{music}}$ not only reflects the mood and theme indicated by the input visuals but also adheres to the stylistic and genre-specific guidelines necessary for generating contextually appropriate music pieces.
-Two type of $Prompt_{bridge}$, for image and video input separately, are shown in Table 2 and 3
-The process can be formulated as below. Given an visual descriptive caption $D_{\text{caption}}$, the module generates a corresponding music-descriptive prompt $D_{\text{music}}$ :
-(4) $D_{\text{music}}=f_{\text{LLM}}(D_{\text{caption}}|Prompt_{bridge})$
-where $f_{\text{LLM}}$ denotes the LLM to transform the descriptive texts into a coherent musical prompt that encapsulates the intended mood, themes, and potentially, the genre of the music to be generated, with the help of $Prompt_{bridge}$.
-Table 2. Prompt template for image-to-music generation.
+Table 5: Prompt template for image-to-music generation.
 Role Content
@@
 assistant Pop dance track with catchy melodies, tropical percussion, and upbeat rhythms, perfect for the beach
-Table 3. Prompt template for video-to-music generation.
+Table 6: Prompt template for video-to-music generation.
 Role Content
@@
 assistant Rock concert with dynamic guitar riffs, precise drumming, and powerful vocals, creating a captivating and electrifying atmosphere, uniting the audience in excitement and musical euphoria.
-By invoking LLMs through API, the model is able to distinguish semantic nuances with high accuracy while ensuring its lightweight nature.
-This capability not only fosters streamlined processing but also facilitates seamless deployment of model services on servers with constrained computational resources.
-3.3. Music Generation Module
-The Music Generation Module utilizes the pre-trained model MusicGen-medium (Copet et al., 2024) to generate music pieces based on the music-descriptive prompts provided by the LLM Understanding & Bridging Module.
-MusicGen is designed to produce high-quality music compositions while accommodating various musical styles and preferences.
-By integrating MusicGen into the Mozart’s Touch framework, we ensure that the generated music aligns closely with the intended mood and theme extracted from the input visuals.
-Given a music-descriptive prompt $D_{\text{music}}$, the Music Generation Module generates a music piece $M$:
-(5) $M=f_{\text{MusicGen}}(D_{\text{music}})$
-where $f_{\text{MusicGen}}$ represents the MusicGen model to transform the music prompt into music composition audio.
-It encapsulates the complex process of interpreting the prompts and translating them into musical elements such as melody, harmony, rhythm, and texture, ensuring that the generated music pieces accurately reflect the intended mood and themes conveyed by the input visuals.
-4. Experiments
-In this section, we assess the image-to-music and video-to-music generation capacities of Mozart’s Touch, with the discussion of two evaluation datasets MUImage and MUVideo, and the evaluation metrics utilized.
-The result of evaluation shows our current state-of-the-art performance in the task of multi-modal music generation.
-4.1. Evaluation Dataset
-To assess our framework’s performance of image-to-music generation, we utilize the MUImage dataset proposed by M2UGen (Hussain et al., 2023a).
-MUImage is assembled by obtaining music samples from the AudioSet (Gemmeke et al., 2017) with corresponding images, which contains 9,966 music-image pairs in total.
-We sampled 2,500 music-image pairs randomly from MUImage as our evaluation dataset.
-For video-to-music generation task, we utilize the MUVideo dataset, which is also proposed by M2UGen.
-We adopted a construction method similar to that of the image-to-music generation task, yielding a corpus of 2,500 music-video pairs for evaluating video-to-music generation task.
-4.2. Evaluation metrics
-For both tasks, we utilize the Frechet Audio Distance (FAD)(Kilgour et al., 2019), Kullback-Leibler divergence (KL) and ImageBind Rank (IB Rank)(Girdhar et al., 2023) as the evaluation metrics.
-FAD is a reference-free evaluation metric for music enhancement algorithms.
-A low score of FAD indicates a high quality of generated music.
-KL scores measure the labels between the original and the generated music.
-When the KL score is low, the generated audios are expected to share similar distributions with the reference music.
-For these two metrics, we utilize the official implementation in PyTorch, where FAD score is supported by the VGGish model.
-IB Rank(Girdhar et al., 2023) is introduced by M2UGen, to assess the alignment between the image/video modality and the generated music.
-Firstly, we use the Image-Bind model to obtain embeddings for the images/videos and the generated music, then calculate their cosine similarity scores and give them a score based on their ranking.
-For IB Rank, High score represents a relatively high ranking among the baselines.
-4.3. Baselines and Details
-For both tasks, we compare Mozart’s Touch with two baselines: CoDi(Tang et al., 2023) and M2UGen(Hussain et al., 2023a).
-We use open-source CoDi model and M2UGen checkpoint files to run inference.
-Our framework runs on one NVIDIA RTX 3090 24GB GPU, and two baselines run on one NVIDIA V100 32GB GPU to load the whole models.
-4.4. Performance Comparison
-Table 4 presents the performance of our framework, Mozart’s Touch, and two baseline models in image-to-music generation.
-The results highlight significant improvements in both the quality and relevance of the music generated by our framework.
-Moreover, Mozart’s Touch surpasses prior state-of-the-art models despite its simpler architecture.
-Table 5 shows the results of video-to-music generation.
-For this task, we observed that Mozart’s Touch still outperforms other models, indicating that our two-step captioning strategy is also highly effective.
-Table 4. Objective comparison of models for image-to-music generation. The best results are made bold.
-Model $FAD_{vgg}$ KL↓ IM Rank↑
-M2UGen 9.166 1.870 0.556
-CoDi 6.674 1.821 0.525
-Mozart’s Touch 4.625 1.169 0.753
-Table 5. Objective comparison of models for video-to-music generation. The best results are made bold.
-Model $FAD_{vgg}$ KL↓ IM Rank↑
-M2UGen 9.047 1.878 0.552
-CoDi 5.055 1.195 0.494
-Mozart’s Touch 4.339 1.048 0.787
-4.5. Subjective Evaluation
-Although we achieve exceptional performance in the objective evaluation, we also believe that quantitative evaluation method has great limitations for music generation tasks.
-The metrics above can effectively measure the quality and relevance of the generated music, but fall short in the understanding of creativity and human feelings, as supported by previous research (Liu et al., 2023b).
-Following previous similar works (Liu et al., 2023b; Kreuk et al., 2023), the generated samples are rated based on i) overall quality (OVL); and ii) relevance to the input image (REL).
-Both OVL and REL metrics have a Likert scale (Likert, 1932) between one and five, where a larger number indicates better performance.
-In this case, We conduct the subjective evaluation involving 125 participants, taking image-to-music generation as example.
-Totally 75 questions are created for the subjective evaluation, which are randomly sampled from our evaluation dataset.
-Each question contains a video with the input image as the visual part and generated (or ground truth) music as the audio.
-20 audios are sampled from ground truth, 20 from M2UGen, 20 from Mozart’s Touch, and 15 from CoDi.
-Each questionnaire comprises ten randomly selected questions.
-Upon subsequent validation by our team, all 75 questions are covered by the total 125 questionnaires.
-Table 6. Subjective comparison of models for image-to-music generation. The best results are made bold.
-Model OVL↑ REL↑
-CoDi 2.95 3.24
-M2UGen 3.77 3.02
-Mozart’s Touch 3.74 3.76
-Ground Truth 3.88 4.08
-The subjective evaluation result is presented in Table 6.
-While our method slightly underperforms in terms of the metrics for overall quality (OVL) when compared to M2UGen, the result shows that there is a notable enhancement in the metric of relevance (REL) to input image, which is consistent with our target to generate corresponding music that aligns the image well.
-4.6. Ablation Studies
-Table 7. Ablation study on image-to-music generation task. The best results are made bold.
-Model $FAD_{vgg}$ KL↓ IM Rank↑
-Mozart’s Touch 4.625  1.170 0.757
-w/o LUBM 3.741 1.121 0.743
-To demonstrate the effectiveness of LLM bridging modality, we conducted a further ablation experiment, comparing the performance of the original system with and without (w/o) the LLM Understanding & Bridging Module (LUBM) in the task of iamge-to-music generation.
-As indicated in the table 7, the framework without LUBM achieves higher scores in the FAD and KL metrics, the two metrics measure the similarity between ground truth and generated audios, rather than the similarity between different modalities.
-On the other side, the framework with LUBM performs better in IB Rank metric.
-This metric utilizes the ImageBind model to encode multi-modal information uniformly, thereby evaluating the similarity between input modality information and generated audio, aligning more closely with the objectives of evaluating multi-modal music generation.
-Therefore, we believe that there is no clear superiority or inferiority between the Mozart’s Touch framework with and without LUBM.
-This once again emphasizes that quantitative evaluation may not always be the best approach for assessing the multi-modal music generation tasks.
-4.7. Case Study
-In this part, we conduct a case study to analyze how our LLM Understanding & Bridging Module (LUBM) mitigates the problem of heterogeneous representations among different modalities.
-By showcasing some representative comparative examples in Figure 2, We demonstrate that the absence of the LUBM does indeed have adverse effects on the generation results.
-The first example illustrates a portrait of Bach.
-Some keywords in the original image description disturb the generation of corresponding music, as they focus on the attributes of image instead of that of music.
-The second example illustrates an anime girl from a visual novel game Atri: My Dear Moments.
-This example shows that insufficiency of music attributions may also mislead the generation of music in a quite different way.
-Figure 2. Case Study
-5. Conclusion
-This paper introduces Mozart’s Touch, a lightweight multi-modal music generation framework that seamlessly integrates LLMs with pre-trained models together.
-Experiments and researches demonstrate the framework’s capability to perform multi-modality understanding and captioning, multi-modal representations bridging, and music generation, resulting in highly aligned music based on the corresponding multi-modal inputs.
-Our future work will aim to refine our prompting strategy for enhanced alignment with multi-modal inputs, conduct further evaluation experiments on our LLM Understanding & Bridging Module, and integrate recent advancements actively into our framework.
-Meanwhile, we will maintain its lightweight characteristic of our framework, ensuring its user-friendliness and expanding its accessibility to a broader scenario.
+Acknowledgements.
+This work was partly sponsored by Undergraduate Training Programs for Innovation and Entrepreneurship of Beijing University of Posts and Telecommunications.
+We sincerely acknowledge that the early stages of this work were inspired by Sylvain Filoni’s prior project at https://huggingface.co/spaces/fffiloni/img-to-music .